OpenStudy (anonymous):

Find the difference 4d^2e - d^2e My guess was 4d^2e

OpenStudy (anonymous):

what about e

OpenStudy (anonymous):

it's 2d^2e that's the answer

OpenStudy (ranga):

4d^2e - d^2e = d^2e( 4 - 1) = 3d^2e